### v1

# Secure Human Oversight of AI: Exploring the Attack Surface of Human Oversight

Abstract
Human oversight of AI is promoted as a safeguard against risks such as inaccurate outputs, system malfunctions, or violations of fundamental rights, and is mandated in regulation like the European AI Act. Yet debates on human oversight have largely focused on its effectiveness, while overlooking a critical dimension: the security of human oversight. We argue that human oversight creates a new attack surface within the safety, security, and accountability architecture of AI operations. Drawing on cybersecurity perspectives, we analyze attack vectors that threaten the requirements of effective human oversight, thereby undermining the safety of AI operations. Such attacks may target the AI system, its communication with oversight personnel, or the personnel themselves. We then outline hardening strategies to mitigate these risks. Our contributions are: (1) introducing a security perspective on human oversight, and (2) providing an overview of attack vectors and hardening strategies to enable secure human oversight of AI.
Keywords Human oversight  $\cdot$ cybersecurity  $\cdot$ attack vectors  $\cdot$ hardening strategies  $\cdot$ AI governance
1 Introduction
Figure 1: The human oversight attack surface encompasses the AI system’s technical infrastructure, the human oversight personnel, and the communication channels between the two sides. Attacks on human oversight aim to undermine oversight personnel’s epistemic access, causal power, self-control, and fitting intentions, the four requirements of effective human oversight [1]. Malicious actors can utilize different attack vectors to undermine one or several of these requirements (see Table 1). In order to protect human oversight from malicious actors, various hardening strategies need to be considered and deployed to counteract the threats posed by the attack vectors (see Table 2). A well-designed combination of hardening strategies can shield against different attack vectors and, thus assure secure human oversight.
Human oversight of artificial intelligence (AI) is often considered a strategy to address risks associated with the use of AI in high-risk contexts [2, 3, 4]. Human oversight personnel are expected to detect inaccurate outputs, outputs that may pose risks to fundamental human rights (e.g., discriminatory outputs), unsafe states of AI systems, and system malfunctions, and to intervene in order to mitigate or minimize these risks [5]. Human oversight is also a key concept in AI governance and ethics [3, 2] and a central requirement for the implementation of high-risk AI systems in emerging AI regulation, such as Article 14 of the European AI Act [6, 3]. However, the idea of relying on humans to oversee AI in order to mitigate operational risks also faces criticism and is expected to encounter substantial challenges in practice [7, 2, 1]. The most prominent criticisms and challenges include doubts about whether humans can effectively oversee AI systems [2], the difficulty of defining what constitutes effective human oversight [1], clarifying who should provide oversight and when it should occur [6], specifying which risks and unwanted system behaviors oversight personnel should detect [5], and determining how to design and evaluate effective human oversight [1].
However, one key challenge that remains underexplored is the security risks that arise from human oversight of AI. Because human oversight will be implemented as part of the safety, security, and accountability architecture of AI systems [3, 7], malicious actors who successfully attack elements related to oversight can significantly undermine the safety of AI operations. In this sense, the human oversight architecture creates a new attack surface for targeting AI operations in high-risk contexts such as critical infrastructure, public administration, or medicine. A security-focused perspective on human oversight is crucial, given the central role oversight already plays, the expectations placed on it in high-risk contexts, and its growing importance as a regulatory requirement in AI governance, such as under the EU AI Act. If we fail to assess and address the security challenges of human oversight, we risk overlooking potential attack vectors in high-risk contexts; and without awareness of these vectors, we will not be able to design effective hardening strategies.
In this paper, we introduce a security-related perspective on human oversight of AI. Specifically, drawing on cybersecurity research, we outline possible attack vectors that threaten the key requirements of effective human oversight (i.e., epistemic access, self-control, causal power, and fitting intentions; [1]). These attacks may target the AI system itself, the communication between the AI system and oversight personnel, or the personnel directly (see Figure 1). We also draw on the cybersecurity literature to discuss hardening strategies that can mitigate the risks posed by such attacks.
This paper makes two key contributions. First, we introduce a security-related perspective on human oversight of AI. This perspective is critically missing from current research, even though human oversight is becoming an integral component of the safety, security, and accountability architecture of AI operations, and thus an increasingly viable target for malicious actors. Second, we provide an overview of possible attack vectors and hardening strategies contextualized for human oversight of AI. We hope this contribution supports AI governance and system design in ensuring not only effective but also secure human oversight of AI.
2 Related Work
2.1 Effective Human Oversight of AI
Research discusses the nature of human oversight of AI, including the what, when, and who of oversight. Regarding the what, many forms of human-AI collaboration in principle involve some degree of human oversight. Across a wide range of human-AI interaction use cases, humans monitor AI system outputs, check their accuracy, request explanations, adapt outputs, intervene when systems malfunction, or take over manual control in the event of failures [8, 9, 10, 11, 12, 13, 14, 15]. In a more targeted sense, and following how emerging research has tentatively defined the concept [1], human oversight specifically refers to humans mitigating risks in AI operations [16, 6]. More concretely, oversight personnel are tasked with monitoring AI systems, detecting failures and malfunctions, identifying unsafe states, flagging inaccurate outputs, and recognizing outputs that may pose risks to fundamental human rights, such as discriminatory outputs [1, 5, 6]. Once such issues are detected, oversight personnel are expected to act in ways that effectively mitigate emerging risks [5, 1]. This can involve taking over manual control, overwriting inaccurate or inadequate outputs [17], or stopping AI operations and triggering intervention by other human stakeholders, such as developers who can adapt the system. Importantly, the AI system under oversight may be operated by another human who is not considered oversight personnel but a user of the system. For example, a patient using an AI chatbot for mental health support [18] would be the user, while oversight personnel monitor the patient–AI interaction to detect risky situations (e.g., if the patient discusses suicidal behavior with the chatbot).
Regarding the when of human oversight, human involvement is important at every stage of AI development: during system design, at run time, and when inspecting errors and failures in hindsight [4, 19, 20]. In a more targeted sense, human oversight usually refers to personnel overseeing AI systems during run time [1, 3]. This can involve dynamic and time-critical situations, such as humans overseeing autonomous vehicles [17, 21], but also less dynamic contexts, such as AI in judicial decision-making or hiring, where there is more time to deliberate about the accuracy or adequacy of AI outputs [5, 22].
Regarding the who of human oversight, any human with the skills necessary to effectively oversee an AI system could serve as oversight personnel [6, 23]. These individuals may be domain experts such as doctors or judges, who can evaluate the adequacy of specific AI outputs based on their expertise. Alternatively, oversight personnel may be trained to detect AI failures without being domain experts in the task performed by the AI. For example, in a medical AI setting, one oversight person might monitor the system’s technical functions on an aggregate level, while medical professionals evaluate individual outputs requiring domain knowledge [23, 5]. Human oversight can also be performed by a single individual or by a team collaboratively overseeing AI systems.
Research has also examined the requirements for effective human oversight. Sterz et al. [1] identify four key aspects: epistemic access, causal power, self-control, and fitting intentions. As we discuss possible attack vectors, we consider how successful attacks might undermine these requirements. Epistemic access means that oversight personnel possess sufficient knowledge about the AI system and understand the consequences of their interventions. Without it, they may fail to detect system failures or inaccurate outputs. Causal power refers to the ability to intervene in ways that actually influence AI processes and outputs. Without causal power, oversight personnel may be unable to overwrite or correct system outputs. Self-control requires oversight personnel to maintain agency over their own actions. For instance, fatigue or intoxication may impair their capacity to fulfill oversight duties. Likewise, being threatened by another person can also undermine self-control, as the actions performed by the oversight personnel are then steered by another agent. Fitting intentions means that personnel are motivated to carry out their responsibilities with the right intentions. In an extreme case, a malicious actor might be hired as oversight personnel and deliberately undermine AI safety. In a less extreme case, poor working conditions may erode motivation and attention, weakening oversight effectiveness [1].
Finally, regarding how to design effective human oversight, insights from the broader field of human-AI collaboration provide valuable guidance (e.g., [8, 24, 16]). Research emphasizes the sociotechnical nature of oversight and highlights the need to design technical, human, and contextual factors that support effective oversight [5, 3, 1]. Technical factors include system interface design, explainability approaches, and interaction modes between humans and AI systems [9, 25, 26, 27]. Human factors include the skills oversight personnel need and the training they receive [23, 16, 5]. Contextual factors include the design of the working environment, time pressure, and the clarity of roles and responsibilities within oversight teams [5, 28, 29].
2.2 Attack Vectors and Hardening Strategies in Cybersecurity
With the implementation of AI systems across diverse fields of application, ensuring the security of deployed solutions has become increasingly urgent. Achieving this requires an in-depth understanding of specific attack vectors and the attack surface within each application domain [30]. Traditionally, an attack vector is defined as a method or pathway that a malicious actor uses to gain access to a network or system. In this paper, we extend the definition to also include methods or pathways that undermine the effectiveness of human oversight. We use the term attack surface to describe the total set of points where a malicious actor could attempt to enter, exploit, or otherwise attack digital and AI infrastructures [30]. For example, research has systematically analyzed attack vectors in autonomous driving systems and defined the corresponding attack surface [31]. Other work has examined the attack surface in federated learning workflows [32]. For human oversight, however, such systematic investigations of the attack surface are still missing.
Relevant related work can be found in the field of AI cybersecurity, which examines attack vectors that explicitly target AI systems. For instance, poisoning attacks manipulate training data to introduce backdoors into AI models that can later be exploited [33, 34, 35]. Adversarial attacks embed imperceptible patterns into inputs to change model behavior, often causing misclassifications or other problematic outcomes [36, 37, 38]. Another line of research targets explainable AI (XAI) methods designed to enhance understanding of AI decisions [25]. Such attacks manipulate explanations to obscure the real reasons behind a model’s decision [39, 40, 41].
Because human oversight extends beyond an isolated AI system, additional relevant research comes from traditional cybersecurity. Distributed denial-of-service (DDoS) attacks flood networks with traffic, making them unresponsive or unavailable [42, 43]. Man-in-the-middle attacks intercept and manipulate data traffic between two endpoints [44]. Another common vector is malware, which – due to its diversity – can be deployed against nearly any digital system [45, 46].
Particularly relevant for defining the attack surface are attacks on the human factor in cybersecurity. Social engineering exploits human vulnerabilities to extract sensitive information or gain access privileges [47, 48]. Two additional but in cybersecurity research less discussed vectors are coercion and bribery, which can likewise be used to gain sensitive information or access privileges [49, 50]. Finally, insider threats exploit individuals inside an organization (e.g., disgruntled employees) to carry out attacks [51]. With the growing threat of cyberattacks, every system that incorporates digital components must be analyzed for potential attack vectors. Our work contributes to this literature by contextualizing attack vectors for the specifics of human oversight of AI.
In response to the many ways digital systems can be attacked, cybersecurity research has developed hardening strategies as methods designed to secure systems against attacks [52, 53, 54]. One line of work establishes methods for securing data traffic in computer networks. Network management systems can detect malicious traffic [55, 56], while encryption protects legitimate communications [57, 58]. To monitor organizational IT systems, intrusion detection systems have been introduced and are continuously improved [59, 60]. Another widely used practice is red teaming, where every system component is rigorously tested by simulating attacks; research continues to optimize red teaming approaches [61, 62]. Beyond technical strategies, organizational measures also play a key role. Transparency of software components is crucial for analyzing systems and identifying potentially harmful functionalities [63, 64]. Other well-established approaches include governance principles and audit frameworks that help secure deployed systems [65, 66]. Finally, like attack vectors, hardening strategies must also address the human factor. Research has explored how to improve system security by training users and raising awareness of risks [67, 68].
In this paper, we extend the literature by contextualizing hardening strategies specifically for securing human oversight of AI against malicious attacks.
3 The Human Oversight Attack Surface
Human oversight can serve as part of the safety, security, and accountability architecture of AI systems, as oversight personnel are expected to detect and mitigate risks. In addition, oversight personnel may be required to actively confirm that an AI system is in a safe state. Human oversight may become even more important in AI operations if it evolves into a central requirement of AI governance and regulation in high-risk contexts [2, 6, 3, 1].
With a more wide-spread implementation, human oversight will also become an increasingly attractive target for malicious actors seeking to corrupt or disrupt AI-based operations. Such actors could be cybercriminals targeting high-risk AI systems for ransom or foreign governments attacking a country’s critical infrastructure in hybrid or open conflict. Furthermore, sufficiently powerful AI agents that are successfully attacked could themselves become malicious actors by attempting to evade oversight [69, 70]. Considering these threats, designing human oversight requires not only ensuring effective oversight [1] but also ensuring secure oversight.
Designing for secure human oversight demands an understanding of its attack surface. The following list of attack vectors inspired by the broader cybersecurity literature is not exhaustive but illustrates the breadth of the human oversight attack surface. It outlines attacks that target the four core requirements of effective human oversight: epistemic access, self-control, causal power, and fitting intentions. Table 1 summarizes these attack vectors and their relation to the requirements for effective human oversight.
Table 1: Attack vectors that target the effectiveness of human oversight. For each attack vector, we indicate the specific requirement for effective human oversight that the attack seeks to undermine.
Attack vectors Affected requirements of effective human oversight
Poisoning attacks Epistemic access, self-control
Adversarial attacks Epistemic access
Explainability attacks Epistemic access
(Distributed) denial of service attacks Causal power
Man-in-the-middle attacks Epistemic access, causal power
Vulnerability exploitation Epistemic access, causal power
Malware Epistemic access, causal power
Social engineering Epistemic access, self-control
Coercion Self-control
Bribery Fitting intentions
Insider threats Epistemic access, causal power, fitting intentions
3.1 Poisoning Attacks
Poisoning attacks target the core functionality of AI models, namely the association of patterns in data with outcomes (e.g., real-valued numbers in regression, class probabilities in classification, or new text/image data in generative AI) [71]. These attacks exploit the fact that modern AI models are often trained on large volumes of (semi-)automatically collected data, for example through web crawlers. Malicious actors can insert triggers, or predefined patterns, into such datasets. When models are trained on compromised data, they may consistently exhibit malicious behaviors whenever a trigger is presented [72].
Poisoning attacks can compromise both the epistemic access and the self-control of oversight personnel. They undermine epistemic access because poisoned samples containing hidden triggers cause AI responses that deviate from normal outputs. A carefully crafted trigger, such as an imperceptible pixel pattern in images [73], can prevent oversight personnel from recognizing the behavioral shift. Even when oversight personnel detect that something is amiss, hidden triggers may still make it difficult to understand the malicious behavior.
Poisoning attacks can also erode self-control. If triggered outputs cause the oversight interface to generate excessive micro-notifications or confirmation requests, malicious actors can exploit this to induce fatigue by overwhelming oversight personnel with trivial but time- and energy-consuming tasks. This not only reduces oversight efficiency by consuming resources but also increases the likelihood of errors once fatigue reaches a critical level, thereby undermining the security of human–AI interaction.
3.2 Adversarial Attacks
Adversarial attacks aim to induce misclassifications in the outputs of an AI model [71]. By feeding the system with adversarial examples, attackers seek either to fool the model or to reduce its accuracy and performance. These adversarial examples typically resemble normal inputs but include subtle perturbations that may not be perceptible to oversight personnel, leading to misclassifications aligned with the attacker’s goals [72].
Adversarial attacks primarily compromise the epistemic access of oversight personnel. A crucial part of understanding AI behavior is knowing how input patterns are mapped to outputs (e.g., which structures or features in an input image correspond to specific output classes). Without this knowledge, reliably determining whether the system behaves acceptably becomes exceedingly difficult. Adversarial perturbations obfuscate the relationship between inputs and outputs, reducing epistemic access to the model’s true behavior. As a result, oversight personnel may fail to detect unwanted model behavior or even system failures.
3.3 Explainability Attacks
Attacks on explainability specifically target explainability and interpretability approaches and tools. These attacks manipulate or deceive the evidence for a model’s reasoning. A key characteristic of such attacks is their ability to alter the explanation while keeping the model’s output unchanged. This means that the AI system may still produce the same output, but the explanation for that output is inaccurate or misleading [74].
Explainability attacks primarily undermine the epistemic access of oversight personnel. Misleading or inaccurate explanations prevent personnel from understanding the AI system’s true behavior. Attackers can exploit these attacks to fabricate behavior that serves their goals. Even with extensive querying, oversight personnel may be unable to detect such manipulations [75]. Moreover, a mismatch between outputs and explanations adversely affects personnel who rely on explanations to identify problematic AI behavior [25, 76].
These attacks can also induce a state of deceptive epistemic access, where the system appears transparent but provides inaccurate explanations. This scenario can be more dangerous than a “black box” system because it fosters unjustified trust and makes detecting incorrect outputs more difficult. Consequently, a key focus of explainability for human oversight must be ensuring the fidelity and robustness of explanations [77, 78].
3.4 (Distributed) Denial of Service Attacks
Denial of service (DoS) attacks refer to methods that render a system inaccessible [79]. The core principle of DoS attacks is to overwhelm a target system with more requests than it can handle, causing the system to become nonfunctional. Various techniques exist to perform a DoS attack, such as exploiting the TCP/IP protocol (SYN flooding), using the ping functionality (ping flooding), or targeting mail servers (mail flooding). A variation of DoS attacks is the distributed denial of service (DDoS) attack, which distributes requests across multiple systems rather than originating from a single source [80].
(Distributed) DoS attacks can undermine the causal power of oversight personnel. Modern computing systems are increasingly implemented as network-based, distributed systems—a standard likely to apply in human oversight settings, where the AI system runs on a cloud-based or on-premise server while oversight personnel interact with it via network protocols. (D)DoS attacks can make either the AI system or the human oversight interface inaccessible. This threat is especially severe for cloud-based AI systems: because the AI system and oversight personnel are physically separated, malicious actors can effectively disable oversight through (D)DoS attacks while the AI system continues to operate in an unsafe state or produces inaccurate outputs.
3.5 Man-in-the-Middle Attacks
An attack vector that can be used in computer networks is the Man-in-the-Middle (MitM) attack [44]. When successful, this strategy allows malicious actors to intercept and control data traffic between two or more systems, granting them access to the exchanged information and enabling them to manipulate it. Additionally, attackers can conceal their presence, tricking the involved systems into believing they are communicating securely with each other [44]. Various methods exist to carry out a MitM attack: ARP spoofing can be used when the attacker and target are on the same local area network (LAN), snarfing can be applied in WLAN-based attacks, and domain name system (DNS) cache poisoning is another possible approach.
MitM attacks can compromise the epistemic access and causal power of oversight personnel, depending on which direction the manipulated information travels. If the attacker intercepts and alters information from the AI system to the oversight personnel, epistemic access is undermined: the true behavior, functionality, or outputs of the AI system may remain hidden.
If the manipulated information flows from the human to the AI system, causal power is compromised. For example, if an attacker modifies commands issued by oversight personnel, they lose the ability to interact effectively with the AI system and cannot intervene in AI processes. This situation is particularly dangerous because the attacker can send manipulated responses back to the personnel, creating an illusion of control and misleading them into believing they are still effectively supervising the AI system.
3.6 Vulnerability Exploitation
Vulnerability exploitation refers to utilizing existing flaws in hardware and software for attacks, such as remote code execution or exfiltration of sensitive data. Since modern digital systems are complex combinations of hardware and software components and their respective supply chains, flaws can accumulate quickly and pose a significant threat. One of the most dangerous vulnerability exploitation is the zero-day exploit [81]. This term refers to the exploitation of a specific vulnerability within a hardware or software component that is unknown even to the manufacturer of the component. This gives the attacker the time advantage of day zero, before the manufacturer can detect, identify and fix the vulnerability.
Attacks based on vulnerability exploitation can target epistemic access or causal power. Attackers can use the vulnerabilities to get unauthorized access to the system and in this way manipulate the system output which the human oversight person uses to detect deviations in system behavior. If the attacker can obfuscate their actions from the human oversight personnel, the personnel are prevented from understanding the real functioning of the AI system. Similar effects can be achieved by manipulating other information that human oversight personnel uses to assess the AI system.
Causal power is affected by vulnerability exploitation if attackers prevent human oversight personnel from appropriately communicating with the AI system. This can be achieved by dropping commands send by the human oversight personnel or disabling crucial components of the human oversight software. Depending on the exploitation, attackers might be able to disguise the attack. For example, and similar to the example for MitM attacks, they could drop commands by human oversight personnel but fake a response by the AI system that suggests the command was successfully executed.
3.7 Malware
Malware is software that executes harmful functionality on an IT system, often without the user’s knowledge. Its potential targets and effects are highly diverse. Ransomware encrypts the file system and typically demands a ransom for decryption. Spyware extracts information from the infected system. These are just two of the most common types, but malware encompasses a wide range of threats designed for various purposes [82, 45].
Due to the diversity of malware functionality, it can compromise epistemic access and causal power of oversight personnel. Regarding epistemic access, malware can alter the core functionality of an AI system. For example, it could insert system prompts into deployed large language models (LLMs), significantly changing their behavior [83]. Malware can also intercept or manipulate data traffic from the AI system to oversight personnel, or modify, encrypt, or obfuscate system logs. These actions prevent personnel from understanding the system’s true functionality and therefore undermine epistemic access.
Causal power can be compromised if malware intercepts or manipulates data traveling from oversight personnel to the AI system. For instance, if malware prevents personnel from interacting with the AI system or locks them out of controlling its core functions, they lose the ability to intervene in system processes or outputs. Similarly, if malware secretly alters commands issued by oversight personnel, it further undermines their causal power.
3.8 Social Engineering
Social engineering refers to attacks that exploit human characteristics [84]. These attacks leverage traits such as fear, respect for authority, or the willingness to help, tricking victims into revealing personal information, disabling security measures, or installing malware. The widespread use of digital communication channels has provided malicious actors with effective and scalable tools for conducting social engineering attacks [85].
Social engineering can compromise epistemic access and self-control of oversight personnel. If an attacker impersonates individuals with technical expertise or authority (e.g., trusted colleagues, management, external experts) they can distort personnel’s understanding of relevant information needed to oversee AI systems. For example, an attacker might pretend to be a technician from the AI provider, convincing human oversight personnel that certain graphical user interface notifications indicate a bug. This could lead personnel to ignore important alerts or install malware disguised as a hotfix. Similarly, attackers impersonating management may convince oversight personnel that specific system notifications are non-critical, framing them as part of corporate processes. In both cases, social engineering feeds false information about the AI system’s functionality to oversight personnel, thereby undermining epistemic access.
Social engineering can also erode self-control. By impersonating senior colleagues or management, attackers can exploit perceived authority to pressure oversight personnel into taking specific actions. If the attack is credible, personnel might follow instructions against their better judgment, believing the directives come from a legitimate authority.
3.9 Coercion
Coercion is a broad concept that extends beyond IT security and involves using force, often through threats or violations, to influence a party. In the context of human oversight, coercion refers to pressure or force that alters the intended actions of oversight personnel.
Coercion undermines the self-control of oversight personnel. By influencing oversight personnel to act according to the instructions of a malicious actor rather than their own judgment, coercion renders human oversight ineffective. Even when oversight personnel detect a need for intervention, they may be unable to act according to their assessment. In extreme cases, a malicious actor could threaten the oversight personnel to gain control over their actions, effectively nullifying the intended benefits of human oversight.
3.10 Bribery
Bribery refers to the act of persuading a person, organization, or authority to act in accordance with the goals of a malicious actor, typically by offering monetary incentives, status, or similar benefits.
Bribery undermines the fitting intentions of human oversight personnel. Usually, the intentions of human oversight personnel should align with fulfilling their oversight duties, such as detecting inaccurate outputs or system failures [1]. A successful bribe can misalign these intentions, causing personnel to act in accordance with the malicious actor’s goals instead. Such goals might include damaging critical infrastructure or compromising the functioning of AI in high-risk applications, such as in hospitals.
3.11 Insider Threats
Insider threats encompass attack scenarios in which malicious actors infiltrate the organization they intend to compromise [86]. Such attacks share the common characteristic that the attacker either has direct access themselves or exploits an employee’s access rights, whether voluntarily or accidentally. Two pathways to insider threats are discussed under the sections on coercion (see 3.9) and bribery (see 3.10). Another common pathway is the exploitation of disgruntled employees.
Insider threats can undermine epistemic access, causal power, and fitting intentions of oversight personnel. They can erode epistemic access by altering or falsifying monitoring data from the AI system. Since human oversight personnel rely on this data to accurately assess AI behavior and intervene when necessary, this prevents them from understanding the system. Insider threats can also obfuscate AI system documentation by deleting or modifying key information, further impeding epistemic access.
Insider threats can also compromise the causal power of oversight personnel. For example, possible sequences of actions that can be taken as a response to detecting problematic system behavior can be ambiguous and the effectiveness of different sequences of actions can differ vastly. Attackers may be able to change organizational oversight policies on how to deal with inaccurate outputs or system malfunctions. In the worst case, such adapted policies can prevent oversight personnel from acting appropriately after detecting problematic system behavior which negatively affects causal power.
Another example of how insider threats may undermine causal power is by achieving and thereby exploit under-staffing. Redirecting resource allocation away from human oversight or preventing human oversight personnel from doing their work (e.g., by force) can reduce the workforce available to human oversight. While this requires substantial access and intervention capabilities, it can prevent oversight personnel from exercising causal power effectively.
Insider threats can additionally undermine the fitting intentions of oversight personnel. Malicious actors with insider access can manipulate the organizational environment to align human oversight personnel’s behavior with their goals. One method is positive punishment, such as establishing a “culture of fear” that discourages human oversight personnel in engaging in behavior that is aligned with their oversight duties. Another method is positive reinforcement, where incentive structures reward behaviors deemed advantageous by the malicious actor.
4 Hardening Human Oversight
Table 2: Hardening strategies are designed to mitigate the risks of (cyber)attacks on human oversight by malicious actors. Each strategy is mapped to the specific attack vectors it protects against.
Hardening strategy Targeted attack vector
Network management (D)DoS attacks
Intrusion Detection System (D)DoS attacks, MitM attacks, malware, insider threats
Encryption MitM attacks
Transparency All attack vectors
Governance principles and audit frameworks Vulnerability Exploitation, insider threats
Red Teaming Poisoning, adversarial, and explainability attacks, vulnerability exploitation
Training of human oversight personnel Social engineering, coercion, bribery
There are various hardening strategies to minimize the risk of malicious actors exploiting attack vectors. The following list of hardening strategies draws on cybersecurity research and practice. Similar to the list of attack vectors, this list is not exhaustive but is intended to inspire approaches for designing secure human oversight. Table 2 provides an overview of mitigation strategies and indicates which attack vectors they can address in the context of human oversight.
4.1 Network Management
As discussed in section 3.4, (D)DoS attacks overwhelm a network by generating excessive traffic. Defending against these attacks requires a multi-step approach [87, 88]. The first step is to deploy monitoring systems that analyze network traffic and trigger alerts if a potential (D)DoS attack is detected. The second step involves methods to distinguish between benevolent traffic (typically generated by humans) and malicious traffic (often produced by bots). Finally, filtering techniques can mitigate (D)DoS attacks by restricting or limiting what users – humans or bots – can do within the network.
Hardening strategy: Establish robust network management procedures within the human oversight architecture.
4.2 Intrusion Detection System
Intrusion detection involves the active monitoring of computers and networks to identify attacks and misuse. An intrusion detection system (IDS) typically combines multiple components and methods to perform this monitoring. The core technical components of an IDS include network or host sensors (depending on the IDS type), a database, a management module, and an analysis module. Host and network sensors monitor the system and collect the necessary information needed for the detection of attacks. The database needs to ensure that the data collected by sensors is stored with a sufficient degree of detail and for a sufficient amount of time. The management module provides all functionalities needed for configuring and calibrating the IDS. The analysis module provides all functionalities necessary for analyzing stored data and performing incident detection. A well-designed and maintained IDS can significantly enhance the cybersecurity of IT systems [89, 59]. Intrusion detection systems may counter multiple attack vectors, including (D)DoS attacks, MitM attacks, malware, and insider threats.
Hardening strategy: Organizations should implement intrusion detection systems to protect their networks broadly and specifically to safeguard human oversight operations and the tools used by oversight personnel from malicious actors.
4.3 Encryption
Encrypted communication is the standard for data traffic on the internet. However, non-encrypted or weakly encrypted data traffic is still commonly used, particularly for communications that do not involve direct interaction between a user and a website. Such vulnerable traffic enables malicious actors to carry out Man-in-the-Middle (MitM) attacks relatively easily, often without requiring technical expertise or expensive software and hardware. To ensure adequate security, modern cryptographic methods with strong security guarantees must be employed to encrypt data traffic [90, 91].
Hardening strategy: All communication between human oversight personnel and the AI system must be encrypted. When personnel and the AI system are not co-located, using a virtual private network to protect data traffic is recommended.
4.4 Transparency
One major challenge in securing AI systems is their complexity. Enhancing transparency of both the system and its communication with human oversight personnel can improve cybersecurity by facilitating the detection of attacks. While transparency itself is not typically considered a cybersecurity measure, it provides access to key security-relevant information, such as implemented components, training data, system functionality, development processes, and dependencies. When combined with other cybersecurity tools, such as vulnerability management, the information made available through transparency increases the likelihood of detecting attacks. Additionally, transparency supports auditing and compliance attestations by making security-relevant information easily accessible and enables model tracking. Overall, ensuring transparency can also help human oversight personnel’s response time, as decision-relevant information becomes more readily available. An appropriate level of transparency can help to counter all of the attack vectors mentioned in this work.
Hardening strategy: Promote transparency to enhance attack detection by making security-relevant information accessible and making potential manipulations of the system easier to identify.
4.5 Governance Principles and Audit Frameworks
Governance principles and audit frameworks are essential for effectively managing security risks in highly automated systems. Governance principles provide the fundamental guidelines, rules, and processes that define how an organization makes decisions and ensures accountability. They establish the framework for achieving strategic goals while minimizing ethical, legal, and operational risks. In the context of AI, these principles specify who is responsible for risk assessment, how transparency is ensured, and which standards apply to data security.
A prominent example of implementing these principles is the AAA Framework, which stands for Prospective Assessments, Audit Trail, and Enforceable Compliance. These three pillars offer a comprehensive approach to strengthening system integrity and human oversight. Prospective Assessments, conducted before system implementation, identify and mitigate potential risks early, incorporating the perspectives of affected individuals and groups [66]. A robust Audit Trail, supported by high-quality data and system transparency, is essential for analyzing errors and assigning responsibility. Enforceable Compliance ensures that legal requirements are actionable, with independent audits verifying compliance and enhancing confidence in system integrity. Another well-established framework is the information security management system (ISMS) specified by ISO/IEC 27001 [92]. An ISMS provides the technical and organizational foundation to ensure integrity and reliability of information such as data, commands, and system statuses. Governance principles and audit frameworks are a countermeasure against vulnerability exploitations and insider threats.
Hardening strategy: Organizations need to establish governance principles and audit frameworks after established standards like the AAA framework and ISO/IEC 27001. These principles and frameworks have to include all components of human oversight that are implemented within an organization.
4.6 Red Teaming
The term red team refers to an independent group that investigates an organization’s IT systems to identify vulnerabilities and security flaws [93]. The red team simulates the behavior of malicious actors, employing all methods that real attackers might use. Its goal is to uncover as many exploitable security issues within the tested system as possible. Red teaming can help counter several different attack vectors, including poisoning attacks, adversarial attacks, explainability attacks, and vulnerability exploitation.
Hardening strategy: Organizations should conduct regular red teaming exercises to detect and address security flaws within their human oversight processes.
4.7 Training of Human Oversight Personnel
Training aims to strengthen the resilience of human oversight personnel against attacks [94]. Training initiatives can help personnel to recognize, resist, and properly respond to manipulation attempts. Training may include social engineering awareness, teaching personnel to identify common manipulation tactics such as phishing or impersonation of authority figures. Practical exercises, such as simulated phishing campaigns or role-play scenarios, reinforce recognition and proper reporting behaviors.
Training can also address resilience against coercion, preparing personnel to respond to threats or undue pressure. This includes protocols for escalating coercion attempts, access to secure whistleblowing channels, and preparation to maintain effective human oversight behavior under stress. Anti-bribery training can teach personnel to recognize undue influence. Clear organizational policies defining conflicts of interest and mandatory disclosure mechanisms further reduce the risk of misaligned intentions. Training of human oversight personal counters social engineering, coercion, and bribery.
Hardening strategy: Organizations should implement recurring, context-specific training programs for human oversight personnel. By institutionalizing such training and embedding it in the organization’s security culture, human oversight personnel can develop the awareness, self-efficacy, and procedural support needed to counteract manipulation attempts, thereby contributing to secure human oversight.
5 Discussion
Research on human oversight of AI has primarily focused on increasing its effectiveness [1, 2, 6, 4]. In this paper, we highlight the importance of complementing this focus with a security-related perspective on human oversight. This perspective is critical because various attack vectors can undermine human oversight effectiveness, a risk that will grow as human oversight becomes more widely implemented as part of the safety, security, and accountability architecture of AI operations. Drawing on cybersecurity research [33, 44], we have introduced various attack vectors targeting human oversight, along with corresponding hardening strategies, to support the development of secure human oversight of AI.
Understanding and designing the technical factors, selecting and training human oversight personnel, and shaping working conditions to foster effective human oversight remains an ongoing challenge for research, practice, and AI governance [1, 2, 95]. Our analyses indicate that this challenge will intensify as human oversight becomes standard practice in high-risk AI use cases. Malicious actors targeting the attack surface of human oversight can compromise the AI system, the communication between humans and AI, or the oversight personnel themselves, aiming to undermine epistemic access, causal power, self-control, and fitting intentions as the core requirements for effective oversight [1]. Consequently, such attacks may increase risks in AI operations by reducing the effectiveness of oversight, disabling oversight entirely, or manipulating oversight personnel’s intentions to exacerbate risks. The attack vector list we present is not exhaustive; future research should explore novel attack vectors specific to human oversight. A potential future scenario involves the deployment of malicious AI agents designed to deceive oversight [70, 69]. Once an AI system models the oversight mechanism, it may attempt to circumvent the restrictions imposed by human oversight, further expanding the attack surface and associated risks.
Given these risks, it is essential to consider how to enhance the security of human oversight. To this end, we presented various hardening strategies drawn from the broader cybersecurity literature [52, 53, 54]. These strategies must be actively applied and adapted to the context of human oversight of AI. Our paper contextualizes these strategies for human oversight, representing a crucial first step toward securing oversight architectures against malicious attacks. Future research needs to evaluate the effectiveness of these hardening strategies and could develop additional strategies tailored specifically to human oversight. Beyond technical hardening, future work could also explore the application of amplified oversight. The core principle of amplified oversight is to employ AI itself to systematically reduce complexity to a level that enables human oversight personnel to fulfill their duties effectively [96]. In future, such approaches to support human oversight may even exceed human capabilities and may become a further viable target for malicious actors to attack.
While this paper focuses on malicious actors targeting human oversight, not all risks emerge from malicious actors. Additional risks can for example arise from counterproductive behavior within organizations [97]. Counterproductive behavior refers to intentional or unintentional actions by employees that go against organizational goal, such as neglecting oversight duties, resisting established procedures, or even deliberately undermining AI operations, which can erode the effectiveness of human oversight. Such behavior may stem from factors like dissatisfaction with working conditions, misaligned incentives, organizational conflict, or lack of recognition [97]. Such counterproductive behavior inside organizations may need closer examination in the future to ensure secure human oversight. Such behavior can undermine the core requirements of effective oversight and may promote active damage to the organization or passive under-performance, thereby introducing risks to human oversight of AI. Although these behaviors are not attacks by malicious actors, they may make attacks by external actors more effective, for example, if an attacker targets a dissatisfied employee. Future works could shed light on those risk factors and suitable hardening strategies such as organizational structures, incentive systems, and thoughtful work design [29]. Well-designed security concepts, policies and work conditions can all together mitigate security risks and prevent counterproductive behavior by ensuring that oversight personnel’s goals remain aligned with organizational objectives.
6 Conclusion
Research on human oversight has primarily focused on its effectiveness – and rightfully so, given the complexity of designing oversight processes that genuinely mitigate risks in AI operations [1, 3, 2]. Adding to this complexity, it will become increasingly important to design for secure human oversight as oversight becomes an integral component of the safety, security, and accountability infrastructure of AI operations. We hope this paper serves as a starting point for research that analyzes the attack surface of human oversight and develops effective hardening strategies to prevent malicious actors from exploiting oversight as a gateway to compromise AI safety.
Acknowledgments
The work of Markus Langer is partially funded by the the Daimler and Benz Foundation as part of the project TITAN – Technologische Intelligenz zur Transformation, Automatisierung und Nutzerorientierung des Justizsystems (grant no. 45-06/24). The work of Kevin Baum is partially funded by the German Federal Ministry of Education and Research (BMBF) as part of the project MAC-MERLin (Grant Agreement No. 01IW24007), and the European Regional Development Fund (ERDF) as well as the German Federal State of Saarland within the scope of the Project ToCERTAIN. In addition, the work of Markus Langer and Kevin Baum is partially funded by the DFG grant 389792660 as part of TRR 248 CPEC — Center for Perspicuous Computing. This work has benefitted from the sixth and last authors’ participation in Dagstuhl Seminar Dagstuhl Seminar 25272 "Challenges of Human Oversight: Achieving Human Control of AI-Based Systems".